# Bayesian Sheaf Neural Networks

### v1

@@ Abstract
Equipping graph neural networks with a convolution operation defined in terms of a cellular sheaf offers advantages for learning expressive representations of heterophilic graph data. The most flexible approach to constructing the sheaf is to learn it as part of the network as a function of the node features. However, this leaves the network potentially overly sensitive to the learned sheaf. As a counter-measure, we propose a variational approach to learning cellular sheaves within sheaf neural networks, yielding an architecture we refer to as a Bayesian sheaf neural network. As part of this work, we define a novel family of reparameterizable probability distributions on the rotation group $SO(n)$ using the Cayley transform. We evaluate the Bayesian sheaf neural network on several graph datasets, and show that our Bayesian sheaf models outperform deterministic sheaf models when training data is limited, and are less sensitive to the choice of hyperparameters.
Keywords: Cellular sheaf, Bayesian neural network, variational inference, Cayley transform
@@ 1 Introduction
Geometric and topological deep learning are rapidly developing research areas which, broadly speaking, aim to use geometric priors about the data of interest to develop neural network architectures using methods and tools from differential geometry, topology, and graph theory [12, 46]. Graph neural networks are a central example of geometric deep learning and have been applied to a wide range of tasks such as recommender systems [54], traffic modeling [42, 55], and predicting properties of chemicals [30]. Building on the basic idea of graph neural networks, architectures such as simplicial convolutional neural networks and hypergraph neural networks extend the graph convolution operation to a higher order convolution operation, which has advantages when the dataset contains salient relationships between groups of nodes rather than just pairs such as citation networks[20, 24], chemical data [8, 7], and neural data [44]. However, such convolution operators suffer from over-smoothing, and thus tend to perform best on homophilic graph data, where nodes connected together by edges or higher dimensional simplices have a high likelihood of belonging to the same class [50, 6].
Several approaches have been taken to address the limited capabilities of GCNs and higher-order variants for heterophilic graph data (i.e. not homophilic) [53, 48, 58, 57]. One approach replaces the graph (or simplicial) convolution operation with a convolution operation defined in terms of a cellular sheaf [32, 6]. A cellular sheaf, briefly, is a data structure defined on a simplicial complex which associates a vector space to each simplex, and a linear transformation known as a restriction map to each pair of incident simplices, that is, a pair of simplices where one simplex is a face of the other. The restriction maps provide additional flexibility in the message passing operation. They allow neighboring nodes to maintain distinct features throughout the layers of the sheaf neural network: an advantageous quality when working with heterophilic graph data.
While sheaf neural networks were developed to address the issue of over-smoothing in graph neural networks, they are still susceptible to many of the issues common to all neural networks, such as overfitting to limited training data and lack of robustness to noisy data, weight initializations, and hyperparameter selections. One approach to mitigating such issues is to apply tools and ideas from Bayesian statistics to deep learning [36, 47]. Bayesian approaches to machine learning have been used to quantify uncertainty of neural network predictions [38, 28], reduce overfitting [27, 29], add robustness to adversarial attacks [9], as well as simply improve the predictions of machine learning algorithms [22, 31, 43].
In this paper, we apply a Bayesian framework to learning cellular sheaves within a sheaf neural network. Specifically, we regard the sheaf Laplacian—an object which generalizes the graph Laplacian to cellular sheaves—as a latent random variable of the sheaf neural network.
@@
    We introduce Bayesian sheaf neural networks (BSNN) in which the sheaf Laplacian operator of the sheaf neural network is a latent random variable whose distribution is parameterized by a function of the input data. Motivated by the evidence lower bound for variational Bayesian inference, a pertinent loss function for our BSNN which contains a Kullback-Leibler (KL) divergence regularization term is derived to efficiently train the BSNN.
@@
    We evaluate BSNN models on several benchmark graph datasets, specifically, on node-classification tasks with limited training data.
@@ 1.1 Related Work
We first discuss existing work on sheaf neural networks. Sheaf neural networks (SNN) were first introduced in [32], where they were shown to outperform graph convolutional networks on a semi-supervised node classification task using a hand-selected cellular sheaf. A more in-depth investigation of sheaf neural networks, sheaf diffusion processes, and the advantage of SNNs for heterophilic graph data was performed in [6], where they also constructed a method of learning a suitable cellular sheaf from the data as part of the sheaf neural network. Variants of sheaf neural networks include SNNs with attention-based edge weights [3], and SNNs defined for hypergraphs rather than cellular complexes [19]. Besides hand-crafting a sheaf for a particular problem or learning a sheaf during training, Barbero et al. [2] construct a sheaf at preprocessing time by assuming the graph data lies on a low dimensional manifold and defining a cellular sheaf in a manner inspired by connections from differential geometry.
@@
Although no existing work has applied a Bayesian framework to sheaf neural networks, we mention some of the ways in which Bayesian neural networks have been combined with geometric machine learning. A variational auto-encoder using graph convolutional layers was considered in [40] for an edge detection task. Davidson et al. [18] discuss variational auto-encoders where the latent variable takes values in a hypersphere. Hajiramezanali et al. [31] combine graph variational autoencoders with graph recurrent networks to outperform existing state-of-the-art methods on dynamic link prediction tasks. Beyond variational autoencoders, Fan et al. model attention weights in graph attention networks as random latent variables, as well as in other architectures which use attention [22]. Given that the inclusion or exclusion of edges in graph datasets can be due to spurious noise or modelling assumptions, Zhang et al. view graph datasets as realizations of parametric families of random graphs, and develop a Bayesian graph convolutional network accordingly [56].
@@
Defining a distribution over cellular sheaves with special orthogonal restriction maps requires a convenient family of distributions over $SO(n)$. The most well-known probability distributions on $SO(n)$ are the matrix Langevin (or von Mises-Fisher) distribution [45, 15] and the matrix Bingham distribution [14]. However such distributions are not easily reparameterizable, which makes incorporating them within neural networks trained via backpropagation difficult. Falorsi et al. [21] use the matrix exponential map to construct reparameterizable distributions on $SO(d)$, however, these distributions in general do not have tractable density functions. In this paper, we define a family of reparameterizable distributions on $SO(n)$ defined via the Cayley transform which we refer to as Cayley distributions. These are distinct from the Cayley distributions defined in [41], and we compare the two families of distributions in Section 4. We also mention [35], where a generalized version of the Cayley transform was used to simulate distributions on Stiefel and Grassmann manifolds.
@@ 1.2 Overview
In Section 2, we review preliminary topics such as cellular sheaves and sheaf Laplacians. We review how the sheaf Laplacian can be used to define a sheaf neural network, and the connection to a sheaf diffusion process. We prove a result regarding the linear separation power of a sheaf diffusion process for sheaves consisting of special orthogonal maps. In Section 3, we describe our variational approach to learning cellular sheaves within a sheaf neural network. In Section 4, we define a reparameterizable family of distributions on $SO(n)$, which are used to define the variational distribution for cellular sheaves with special orthogonal restriction maps. In Section 5, we evaluate our BSNN on several node classification tasks for heterophilic web-page datasets. We end with a final discussion, as well as potential directions for future work in Section 6.
@@ 2 Preliminaries
2.1 Cellular Sheaves
A $k$--simplex is the convex hull of $k+1$ affinely independent points in $\mathbb{R}^{n}$. Familiar examples include points, lines, triangles, and tetrahedrons, which are 00, $1$, $2$, and $3$--simplices respectively. A face $\tau$ of a simplex $\sigma$ is the simplex spanned by any subset of the points defining $\sigma$. In this case we denote $\tau\trianglelefteq\sigma$, and may say that $\tau$ is incident with $\sigma$. A simplicial complex $K$ is a collection of simplices in $\mathbb{R}^{n}$ such that for any simplex in $K$, all of its faces are also in $K$, and whenever two simplices $\sigma,\tau$ in $K$ have a non-empty intersection $\sigma\cap\tau$, their intersection is a face of both $\sigma$ and $\tau$.
A cellular sheaf $\mathcal{F}$ on a simplicial complex $K$ is a structure which assigns a vector space $\mathcal{F}(\sigma)$ to each simplex $\sigma\in K$, and assigns to each incident pair $\tau\trianglelefteq\sigma$ a linear transformation $\mathcal{F}_{\tau\trianglelefteq\sigma}:\mathcal{F}(\tau)\to\mathcal{F}(\sigma)$ such that $\mathcal{F}_{\sigma\trianglelefteq\sigma}:\mathcal{F}(\sigma)\to\mathcal{F}(\sigma)$ is the identity map for every $\sigma$, and with the compatibility condition that $\mathcal{F}_{\tau\trianglelefteq\rho}=\mathcal{F}_{\sigma\trianglelefteq\rho}\circ\mathcal{F}_{\tau\trianglelefteq\sigma}$ whenever $\tau\trianglelefteq\sigma\trianglelefteq\rho$. In keeping with the terminology from sheaf theory [11], the vector space $\mathcal{F}(\sigma)$ is referred to as the stalk of $\mathcal{F}$ over $\sigma$, and the transformations $\mathcal{F}_{\tau\trianglelefteq\sigma}$ are referred to as restriction maps. Cellular sheaves can be defined on cell complexes, a slight generalization of simplicial complexes, hence their name. For simplicity of presentation, we restrict our attention to simplicial complexes.
The motivation for considering cellular sheaves in the context of geometric deep learning is that given a cellular sheaf $\mathcal{F}$, there is a corresponding sheaf Laplacian $L_{\mathcal{F}}$. Just as how the message passing operation in many graph convolutional neural networks is defined in terms of the graph Laplacian $L=D-A$ [40], the sheaf Laplacian $L_{\mathcal{F}}$ can be used to define a message passing operation.
To begin defining the sheaf Laplacian for a cellular sheaf $\mathcal{F}$ over a finite simplicial complex $K$, assume a total order is placed on the vertices of $K$, so that each simplex $\sigma\in K$ may be written $\sigma=[v_{0},v_{1},\dots,v_{n}]$ with $v_{0}<v_{1}<\cdots<v_{n}$. Given a $k$--simplex $\sigma=[v_{0},v_{1},\dots,v_{k}]$, we denote its $(k-1)$--dimensional faces by $\partial_{i}(\sigma):=[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$ for $i\in\{0,\dots,k\}$, where $\hat{v}_{i}$ denotes the omission of vertex $v_{i}$.
An $n$--cochain of $(K,\mathcal{F})$ is an element $x\in C^{n}(K;\mathcal{F})$, where $C^{n}(K;\mathcal{F}):=\bigoplus_{\sigma\in K_{n}}\mathcal{F}(\sigma)$ denotes the direct sum of the vector spaces $\mathcal{F}(\sigma)$ over all $n$--dimensional simplices $\sigma\in K_{n}$. Given a cochain $x\in C^{n}(K;\mathcal{F})$ and an $n$--simplex $\sigma$, let $x_{\sigma}$ denote the value of $x$ at $\mathcal{F}(\sigma)$, i.e. $x_{\sigma}=\pi_{\sigma}(x)$ for the projection $\pi_{\sigma}:C^{n}(K;\mathcal{F})\to\mathcal{F}(\sigma)$. The $n$--th coboundary map $\delta_{n}:C^{n}(K;\mathcal{F})\to C^{n+1}(K;\mathcal{F})$ is the linear transformation given by defining the value of $\delta_{n}(x)$ at an $(n+1)$--simplex $\sigma$ to be
$\delta_{n}(x)_{\sigma}=\sum_{i=0}^{n+1}(-1)^{i}\mathcal{F}_{\partial_{i}\sigma\trianglelefteq\sigma}((x)_{\partial_{i}\sigma}).$ (1)
The $n$--th sheaf Laplacian of $(K,\mathcal{F})$ is the linear transformation $L_{\mathcal{F}}^{n}:C^{n}(K;\mathcal{F})\to C^{n}(K;\mathcal{F})$ defined by
$L^{n}_{\mathcal{F}}=\delta_{n-1}\delta_{n-1}^{*}+\delta_{n}^{*}\delta_{n},$ (2)
where $\delta_{n}^{*}:C^{n+1}(K;\mathcal{F})\to C^{n}(K;\mathcal{F})$ denotes the adjoint of $\delta_{n}$. In the case that $\mathcal{F}(\sigma)=\mathbb{R}$ for all $\sigma\in K$ and $\mathcal{F}_{\tau\trianglelefteq\sigma}$ is the identity map on $\mathbb{R}$ for all pairs $\tau\trianglelefteq\sigma$, the $n$--th sheaf Laplacian $L_{\mathcal{F}}^{n}$ is simply the combinatorial Laplacian of the simplicial complex $K$ [25]. For further information on cellular sheaves and sheaf Laplacians, we refer the interested reader to [33].
2.1.1 Cellular Sheaves over Graphs
Although the sheaf Laplacians $L_{\mathcal{F}}^{n}$ for multiple dimensions $n$ could be used as message passing operations within a sheaf neural network, due to the prevalence of graph-based data, one is primarily concerned with cellular sheaves over graphs, which are $1$--dimensional simplicial complexes, and the 00th sheaf Laplacian $L_{\mathcal{F}}^{0}$, denoted by $L_{\mathcal{F}}$.
Given a sheaf $\mathcal{F}$ over a graph $G$ with $n$ nodes, and assuming the stalk $\mathcal{F}(u)$ over each node $u\in V(G)$ has the same dimension $d$, the sheaf Laplacian, $L_{\mathcal{F}},$ may be represented as a block matrix of size $nd\times nd$. The off-diagonal blocks $L_{ij}\in\mathbb{R}^{d\times d}$ are $L_{ij}=-\mathcal{F}_{u_{j}\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{u_{i}\trianglelefteq e},$ whenever there exists an edge between the nodes $u_{i}$ and $u_{j}$, and $L_{ij}=0$ otherwise. The diagonal blocks are $L_{ii}=\sum_{u_{i}\trianglelefteq e}\mathcal{F}_{u_{i}\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{u_{i}\trianglelefteq e},$ where the sum is over all edges $e$ such that $u_{i}\trianglelefteq e$. If we let $D$ denote the $nd\times nd$ block diagonal matrix whose diagonal blocks are $L_{ii}$, and all other blocks are zero matrices, then the normalized 00th sheaf Laplacian $\Delta_{\mathcal{F}}$ defined in [32] is $\Delta_{\mathcal{F}}:=D^{-1/2}L_{\mathcal{F}}D^{-1/2}.$ Note $D$ is symmetric positive semi-definite, hence there exists a matrix $D^{1/2}$ such that $D^{1/2}D^{1/2}=D$. Moreover, we only consider sheaves such that the restriction maps $\mathcal{F}_{u\trianglelefteq e}$ are invertible, in which case $D$ is invertible, and so $D^{-1/2}$ is well-defined.
@@ 2.2 Sheaf Diffusion
Let $G$ be a graph with $n$ nodes and with $d$--dimensional node features, and let $\mathcal{F}$ be a cellular sheaf over $G$ with $d$--dimensional stalks. We represent the node features as an $nd$--dimensional vector $X\in\mathbb{R}^{nd}$ so that the normalized sheaf Laplacian $\Delta_{\mathcal{F}}$, as a $nd\times nd$ block matrix, updates $X$ by matrix-vector multiplication $\Delta_{\mathcal{F}}X$. Motivated by how the graph convolutional layers of [40] can be viewed as a discretization of a heat diffusion process $\tfrac{d}{dt}X(t)=-\Delta X(t)$ defined in terms of the normalized graph Laplacian $\Delta$, the authors of [6] consider a similar diffusion process which now uses the sheaf Laplacian:
@@
Here $W_{t,1}\in\mathbb{R}^{d\times d}$ and $W_{t,2}\in\mathbb{R}^{f\times f}$ are weight matrices for layer $t$, $\sigma$ is a non-linear activation function, $\otimes$ denotes a Kronecker product.
@@
As a measure of the expressive power of sheaf layers based on Eq. (4), Bodnar et al. [6] study the ability of the sheaf diffusion process in Eq. (3) to linearly separate the features of nodes belonging to different classes.
Definition 2.1.
@@
A class of sheaves $\mathcal{H}^{d}$ with $d$--dimensional stalks is said to have linear separation power over a class of labeled graphs $\mathcal{G}$ if for every $G\in\mathcal{G}$, there exists a sheaf $\mathcal{F}\in\mathcal{H}^{d}$ such that the sheaf diffusion process determined by $\mathcal{F}$ on $G$ linearly separates the features $X(t)$ of the classes of $G$ in the time limit $t\to\infty$, for almost all initial conditions $X(0)$.
@@
As part of a sheaf learning mechanism within a sheaf neural network, a fixed class of sheaves $\mathcal{H}^{d}$ is parameterized, and the input data is mapped to a suitable sheaf $\mathcal{F}\in\mathcal{H}^{d}$ through this parameterization. For a node classification task with $C$ classes, the collection of sheaves $\mathcal{H}^{d}$ used would ideally have linear separation power over the collection of connected graphs with $C$ classes. While one could always take $\mathcal{H}^{d}$ to be as general as possible, it can be preferable to use a more restricted collection of sheaves, as this can reduce the number of model parameters involved in learning a cellular sheaf. For example, learning cellular sheaves whose restriction maps are diagonal or special orthogonal matrices involves fewer parameters than when the restriction maps are general linear maps. Recall that a special orthogonal matrix, i.e. element of $SO(n)$, is a matrix $Q$ such that $Q^{T}=Q^{-1}$ and $\det(Q)=1$.
@@
It was shown that the class of sheaves $\mathcal{H}^{d}_{\operatorname{diag}}$ with invertible diagonal restriction maps has linear separation power over the collection $\mathcal{G}_{C}$ of connected graphs with $C$ classes whenever $C\leq d$ [6]. However, for $\mathcal{H}_{so}^{d}$, the class of sheaves with $d\times d$ special orthogonal restriction maps, [6, Proposition 13] only guarantees the existence of a $d$ such that $\mathcal{H}_{so}^{d}$ has linear separation power over $\mathcal{G}_{C}$ when $C\leq 8$.
Remark 2.2.

### v2

@@ Abstract
Equipping graph neural networks with a convolution operation defined in terms of a cellular sheaf offers advantages for learning expressive representations of heterophilic graph data. The most flexible approach to constructing the sheaf is to learn it as part of the network as a function of the node features. However, this leaves the network potentially overly sensitive to the learned sheaf. As a counter-measure, we propose a variational approach to learning cellular sheaves within sheaf neural networks, yielding an architecture we refer to as a Bayesian sheaf neural network. As part of this work, we define a novel family of reparameterizable probability distributions on the rotation group $SO(n)$ using the Cayley transform. We evaluate the Bayesian sheaf neural network on several graph datasets, and show that our Bayesian sheaf models achieve leading performance compared to baseline models and are less sensitive to the choice of hyperparameters under limited training data settings.
Keywords: Cellular sheaf, Bayesian neural network, variational inference, Cayley transform, uncertainty quantification.
@@ 1 Introduction
Geometric and topological deep learning are rapidly developing research areas which, broadly speaking, aim to use geometric priors about the data of interest to develop neural network architectures using methods and tools from differential geometry, topology, and graph theory [14, 55]. Graph neural networks are a central method in geometric deep learning and have been applied to a wide range of tasks such as recommender systems [64], traffic modeling [51, 65], and predicting properties of chemicals [35]. Building on the basic idea of graph neural networks, architectures such as simplicial and hypergraph neural networks extend the graph convolutions to higher-order operations, which have advantages when the dataset contains salient relationships among groups of nodes rather than just node pairs, as in citation networks [26, 30], chemical data [9, 8], and neural data [53]. However, such convolution operators suffer from over-smoothing, and thus tend to perform best on homophilic graph data, where connected nodes have a high likelihood of belonging to the same class [59, 10].
Several approaches have been taken to address the limited capabilities of GCNs and higher-order variants for heterophilic graph data (i.e., not homophilic) [63, 57, 69, 67]. One approach replaces the graph convolution operation with one defined in terms of a cellular sheaf [40, 10]. A cellular sheaf on a graph assigns a vector space to each vertex and each edge, along with linear maps, known as restriction maps, that relate the data on edges to the data on their incident vertices. Restriction maps provide additional flexibility in the message passing operation. They allow neighboring nodes to maintain distinct features throughout the layers of the sheaf neural network: an advantageous quality when working with heterophilic graph data.
While sheaf neural networks were developed to address the issue of over-smoothing in graph neural networks, they are still susceptible to many of the issues common to all neural networks, such as overfitting to limited training data and lack of robustness to noisy data, weight initializations, and hyperparameter selections. An approach to mitigating such issues is to apply tools and ideas from Bayesian statistics to deep learning [44, 56]. Bayesian approaches to machine learning have been used to quantify the uncertainty in neural network predictions [46, 33], reduce overfitting [32, 34], add robustness to adversarial attacks [11], and to improve the overall predictive performance of machine learning algorithms [28, 37, 52].
In this paper, we apply a Bayesian framework to learning cellular sheaves within a sheaf neural network. Specifically, we consider the sheaf Laplacian as an object that enables the integration of the graph Laplacian with cellular sheaves, treating it as a latent random variable of the sheaf neural network.
@@
    We introduce Bayesian sheaf neural networks (BSNN) in which the sheaf Laplacian operator of the sheaf neural network is a latent random variable whose distribution is parameterized by a function of the input data. Motivated by the evidence lower bound for variational Bayesian inference, we derive a pertinent loss function for our BSNN which contains a Kullback-Leibler (KL) divergence regularization term, enabling efficient training.
@@
    We incorporate uncertainty quantification to rigorously evaluate our BSNN models on several benchmark graph datasets for node classification tasks, presenting extensive numerical results and comprehensive sensitivity analyses to validate robustness and performance.
@@ 1.1 Related Work
We first discuss existing work on sheaf neural networks. Sheaf neural networks (SNN) were first introduced by [40], where they were shown to outperform graph convolutional networks on a semi-supervised node classification task using a hand-selected cellular sheaf. A more in-depth investigation of sheaf neural networks, sheaf diffusion processes, and the advantage of SNNs for heterophilic graph data was performed in [10], where they also constructed a method of learning a suitable cellular sheaf from the data as part of the sheaf neural network.
Building on these foundational approaches, several variants of SNNs have been proposed that extend their applicability or enhance their expressiveness. They include SNNs with attention-based edge weights [4] and SNNs defined for hypergraphs rather than cellular complexes [25]. In addition to hand-crafting a sheaf for a particular problem or learning a sheaf during training, [3] construct a sheaf at preprocessing time by assuming the graph data lie on a low-dimensional manifold and defining a cellular sheaf in a manner inspired by connections from differential geometry.
@@
While SNNs have not yet been explored through a Bayesian lens, uncertainty modeling and probabilistic reasoning have played an increasing role in geometric deep learning more broadly. A variational auto-encoder using graph convolutional layers was considered in [48] for an edge detection task. [23] discuss variational auto-encoders where the latent variable takes values in a hypersphere. [37] combine graph variational autoencoders with graph recurrent networks to outperform existing state-of-the-art methods on dynamic link prediction tasks. Beyond variational autoencoders, [28] model attention weights in graph attention networks as random latent variables, as well as in other architectures which use attention. Given that the inclusion or exclusion of edges in graph datasets can be due to spurious noise or modeling assumptions, [66] view graph datasets as realizations of parametric families of random graphs, and develop a Bayesian graph convolutional network accordingly.
@@
A key component of a Bayesian formulation for sheaf neural networks is defining a distribution over the space of restriction maps, which often lie in the special orthogonal $SO(n)$. The most well-known probability distributions on $SO(n)$ are the matrix Langevin (or von Mises-Fisher) distribution [54, 19] and the matrix Bingham distribution [18]. However, such distributions are not easily reparameterizable, which makes incorporating them within neural networks trained via backpropagation difficult. While the matrix exponential used by [27] provides a reparameterizable mapping to $SO(d)$, it lacks tractable densities.
In this paper, we define a family of reparameterizable distributions on $SO(n)$ using the Cayley transform which we refer to as Cayley distributions. These are distinct from the Cayley distributions defined in [50], and we compare the two families of distributions in Section 3.3. Moreover, [43] used a generalized version of the Cayley transform to simulate distributions on Stiefel and Grassmann manifolds, illustrating the broader applicability of Cayley-type methods in geometric statistics.
@@ 1.2 Overview
In Section 2, we review preliminary topics such as cellular sheaves and sheaf Laplacians. We review how the sheaf Laplacian can be used to define a sheaf neural network, and the connection to a sheaf diffusion process. We also review the Cayley distribution, variational inference, and several uncertainty quantification metrics. In Section 3, we describe our variational approach to learning cellular sheaves in a sheaf neural network. We prove a result regarding the linear separation power of a sheaf diffusion process for sheaves consisting of special orthogonal maps. We define a reparameterizable family of distributions on $SO(n)$, which is used to define the variational distribution for cellular sheaves with special orthogonal restriction maps. In Section 4, we evaluate our BSNN models on several node classification tasks and measure their predictive uncertainty. We end with a final discussion, as well as potential directions for future work in Section 5.
@@ 2 Preliminaries
2.1 Cellular Sheaves over Graphs
Definition 1.
A cellular sheaf $\mathcal{F}$ on an undirected graph $G=(V,E)$ assigns a vector space $\mathcal{F}(v)$ to each node $v\in V$, a vector space $\mathcal{F}(e)$ to each edge $e\in E$, and a linear map $\mathcal{F}_{v\trianglelefteq e}:\mathcal{F}(v)\to\mathcal{F}(e)$ for each incident pair $v\trianglelefteq e$ (i.e., node $v$ is incident to edge $e$).
Remark 1.
In the language of sheaf theory, $\mathcal{F}(v)$ is called the stalk over $v$, and $\mathcal{F}_{v\trianglelefteq e}$ is called a restriction map [13].
Definition 2.
The spaces of 0-cochains and $1$-cochains are defined as the direct sums
$C^{0}(G;\mathcal{F}):=\bigoplus_{v\in V}\mathcal{F}(v)\quad\text{and}\quad C^{1}(G;\mathcal{F}):=\bigoplus_{e\in E}\mathcal{F}(e),$
corresponding to the vector spaces assigned to the nodes and edges, respectively.
Definition 3.
Let $x\in C^{0}(G;\mathcal{F})$, $x_{v}$ denotes the value of $x$ at $\mathcal{F}(v)$. The coboundary map $\delta:C^{0}(G;\mathcal{F})\to C^{1}(G;\mathcal{F})$ is defined componentwise by
$(\delta x)_{e}=\mathcal{F}_{v\trianglelefteq e}(x_{v})-\mathcal{F}_{u\trianglelefteq e}(x_{u}),$
where $e$ is an edge oriented from $u$ to $v$ (this orientation is chosen arbitrarily to define a consistent sign convention; the graph itself remains undirected).
Definition 4.
The sheaf Laplacian $L_{\mathcal{F}}:C^{0}(G;\mathcal{F})\to C^{0}(G;\mathcal{F})$ is defined as
$L_{\mathcal{F}}:=\delta^{\top}\delta.$
[…]
Remark 2.
Analogously to how message passing in graph convolutional networks uses the graph Laplacian L=DAL=D-A [48], the sheaf Laplacian $L_{\mathcal{F}}$ defines message passing on cellular sheaves. For further information on cellular sheaves and sheaf Laplacians, we refer the interested reader to [39].
Definition 5.
The normalized 0th sheaf Laplacian $\Delta_{\mathcal{F}}$ associated with a cellular sheaf $\mathcal{F}$ over a graph $G$ is defined as
$\Delta_{\mathcal{F}}:=D^{-1/2}L_{\mathcal{F}}D^{-1/2},$
where $D$ is the block-diagonal matrix whose diagonal blocks are the $L_{ii}$ and whose off-diagonal blocks are zero matrices [40].
Remark 3.
The matrix $D$ is symmetric positive semi-definite, and therefore there exists a matrix $D^{1/2}$ such that $D^{1/2}D^{1/2}=D$. Moreover, we only consider restriction maps $\mathcal{F}_{u\trianglelefteq e}$ that are invertible, ensuring $D$ is invertible and $D^{-1/2}$ is well-defined.
@@ 2.2 Sheaf Diffusion
Let $G$ be a graph with $n$ nodes and with $d$-dimensional node features, and let $\mathcal{F}$ be a cellular sheaf over $G$ with $d$-dimensional stalks. We represent the node features as an ndnd-dimensional vector $X\in\mathbb{R}^{nd}$ so that the normalized sheaf Laplacian $\Delta_{\mathcal{F}}$, as an $nd\times nd$ block matrix, updates $X$ by matrix-vector multiplication $\Delta_{\mathcal{F}}X$. Motivated by how the graph convolutional layers of [48] can be viewed as a discretization of a heat diffusion process $\tfrac{d}{dt}X(t)=-\Delta X(t)$ defined in terms of the normalized graph Laplacian $\Delta$, a similar diffusion process uses the sheaf Laplacian:
@@
Here $W_{t,1}\in\mathbb{R}^{d\times d}$ and $W_{t,2}\in\mathbb{R}^{f\times f}$ are weight matrices for layer $t$, $\sigma$ is a nonlinear activation function, $\otimes$ denotes a Kronecker product.
@@
As a measure of the expressive power of the sheaf layers based on Eq.(2), [10] study the ability of the sheaf diffusion process in Eq.(1) to linearly separate the features of nodes belonging to different classes.
Definition 6.
@@
A class of sheaves $\mathcal{H}^{d}$ with $d$-dimensional stalks is said to have linear separation power over a class of labeled graphs $\mathcal{G}$ if for every $G\in\mathcal{G}$, there exists a sheaf $\mathcal{F}\in\mathcal{H}^{d}$ such that the sheaf diffusion process determined by $\mathcal{F}$ on $G$ linearly separates the features $X(t)$ of the classes of $G$ in the time limit $t\to\infty$, for almost all initial conditions $X(0)$.
As part of a sheaf learning mechanism within a sheaf neural network, a fixed class of sheaves $\mathcal{H}^{d}$ is parameterized, and the input data is mapped to a suitable sheaf $\mathcal{F}\in\mathcal{H}^{d}$ through this parameterization.
@@
For a node classification task with $C$ classes, the collection of sheaves $\mathcal{H}^{d}$ should ideally have linear separation power over the collection of connected graphs with $C$ classes. While one could always take $\mathcal{H}^{d}$ to be as general as possible, learning cellular sheaves whose restriction maps are diagonal or special orthogonal matrices involves fewer parameters. Recall that a special orthogonal matrix, that is, an element of $SO(n)$, is a matrix $Q$ such that $Q^{T}=Q^{-1}$ and $\det(Q)=1$.
@@
The class of sheaves $\mathcal{H}^{d}_{\operatorname{diag}}$ with invertible diagonal restriction maps was shown to have linear separation power over the collection $\mathcal{G}_{C}$ of connected graphs with $C$ classes whenever $C\leq d$ [10]. However, for $\mathcal{H}_{so}^{d}$, the class of sheaves with $d\times d$ special orthogonal restriction maps, as discussed in Proposition 13 of [10], only guarantees the existence of a $d$ such that $\mathcal{H}_{so}^{d}$ has linear separation power over $\mathcal{G}_{C}$ when $C\leq 8$.
Remark 4.